19. Graph the inequality 3x + 2y < 6.
Answer #1

ANSWER:

Answer: Given: 3x4 24 26 First • get pet it in y=mate format So, 3x+2y < 24 s 6-3x 24 <-3x+6 y <-3x+6 2 Y 2 - 2x + 2 y < (-2x

3x+2y=6 6. 3(0) +2y=6 2y=6 Y- = 3 y=3 point (0,3) in y=0 putting equation (1) 3ut2y = 6 3x+2(0) = 6 3x = 6 x = 6 = 2 point (2

3x + 2y < 6 2y < 6-30 y < 6-3x 2 y < 3 - 3 x 2 points (0,3) (2,0) Y 6 5 3 2 х -5 -4 -3 -2 -1 1 2 3 4 5 6 -1 -2 -3 -4 -5